OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? x+y = -1 x-3y = 11 A. (2, −3) B. (−2, 1) C. (5, −2) D. (−6, 5)

OpenStudy (anonymous):

plz help i will medal! and fan!

OpenStudy (anonymous):

i think it is a

OpenStudy (calculusxy):

\[\color \green{x + y = -1}\] \[\color \green{x - 3y = 11}\] \[y = -1 - x \] \[x - 3(-1 - x) = 11 \rightarrow x + 3 + 3x = 11 \rightarrow 4x + 3 = 11 \rightarrow \color \red{x = 2}\] \[\color \purple{2 - 3y = 11} \rightarrow -3y = 9 \rightarrow y= -3\] \[\large \tt \color \green{(2, -3)}\]

OpenStudy (anonymous):

omg thank you so much! here is a medal!

OpenStudy (calculusxy):

You're welcome :)

OpenStudy (anonymous):

i am toats taking notes on this!

OpenStudy (calculusxy):

I am happy that this seems helpful to you!

OpenStudy (anonymous):

it is thx i have a couple more like this but i will work them out and u check i will tag u when i post thx again!

OpenStudy (calculusxy):

Well, I have to go now but I might come back on OS.

OpenStudy (anonymous):

kk bye!